# On multi-graded Proj schemes

### v3

@@ 1.3 Contents
We now present the results contained in this document. Let as above $M$ be an abelian group and $A$ be a commutative ring endowed with a grading by $M$. In Sections 2 and 3 we define the notions of $M$-relevant families of $A$ (following Brenner–Schröer, see Definition 2.3) and of potions associated with these families, and explain the definition of the Proj scheme ${\operatorname{Proj}}^{M}(A)$ of $A$. The potion associated with a relevant family $(f_{i}:i\in I)$ is the degree-zero part in the localization of $A$ with respect to the multiplicative subset generated by $(f_{i}:i\in I)$; see Definition 2.9. The spectra ${\operatorname{Spec}}(A_{(S)})$ are called potion schemes and are denoted $D_{\dagger}(S)$ (cf. Construction 3.1), and the scheme ${\operatorname{Proj}}^{M}(A)$ is defined by glueing the potions schemes $D_{\dagger}(S)$ over all relevant families (cf. Construction 3.1). This glueing is possible by the magic of potions (cf. Proposition 2.10). In other words, Section 2 focuses on commutative algebra around potions while Section 3 applies this material to explain the construction of Proj schemes.
In Sections 2 and 3, we also prove some new basic results. In particular, using the concept of quasi-relevant element (Definition 2.11), we prove that the Proj scheme of a tensor product of graded rings (with the product grading) is the fiber product of the Proj schemes of the rings (Proposition 3.15). We prove some compatibility results regarding the radical of the ideal generated by relevant elements (Propositions 2.5 and 2.7). In Example 2.13, we explain the relation between potions and dilatations of rings (as studied in [MRR23, Ma23b]). We also show that the Proj construction is functorial (Proposition 3.12). In §3.5 we discuss multi-centered blowups and its relations with dilatations.
In Section 4 we explain how to realize flag varieties of reductive groups, and some vector bundles attached to representations of a Borel subgroup, as the Proj scheme of a natural graded ring (without any choice of strictly dominant character). Finally, in Section 5 we study, given an $M$-graded commutative ring $A$, the natural functors relating the categories of $M$-graded $A$-modules and quasi-coherent sheaves on ${\operatorname{Proj}}^{M}(A)$. In particular, we explain how, under a certain technical assumption, one can describe $\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$ as a Serre quotient of the category of $M$-graded $A$-modules, and therefore obtain a version of Serre’s celebrated theorem.
Many of these results are counterparts of well-known results regarding Grothendieck’s Proj construction, which can be found e.g. in [StP].
@@ 1.6.1 Rings
All the rings we will consider in the paper will be tacitly assumed to be unital. (In practice, all the rings we want to consider will be commutative, but we will mention this assumption when it is necessary.)
@@ 1.6.2 Graded rings
Let us make explicit our conventions on graded rings. For this we fix an arbitrary abelian group $M$.
Definition 1.1.
An $M$-graded ring is a ring $A$ endowed with a direct sum decomposition $A=\bigoplus_{m\in M}A_{m}$ such that $A_{m}\cdot A_{m^{\prime}}\subset A_{m+m^{\prime}}$ for all $m,m^{\prime}\in M$.
Definition 1.2.
Let $A$ be an $M$-graded ring. An element $a\in A$ is called homogeneous if it belongs to $A_{m}$ for some $m\in M$. In this case, if $a$ is nonzero, then the element $m$ is unique and denoted $\deg(a)$ and called the degree of $a$.
@@ 1.6.3 Other conventions
If $A$ is a ring and $a\in A$ is an element, we set $a^{\mathbb{N}}:=\{a^{n}:n\in{\mathbb{N}}\}\subset A$. For any $a\in A$, by convention we have $a^{0}=1$.
@@
If $N$ is any commutative monoid, there is a universal homomorphism from $N$ to a group $N^{\mathrm{gp}}$ (cf. [Og18, §1.3]). If $N$ is a submonoid of a given abelian group $M$, then the group $N^{\mathrm{gp}}$ naturally identifies with the subgroup $N-N:=\{n-n^{\prime}:n,n^{\prime}\in N\}$ of $M$.
@@ 2.2 Relevant families
Let $S$ be a homogeneous multiplicative subset of $A$. We denote by $\deg(S)$ the subset of $M$ consisting of degrees of nonzero homogeneous elements of $S$ (cf. Definition 1.2).
@@
    A homogeneous multiplicative subset $S$ of $A$ is called $M$-relevant (or just relevant if $M$ is clear from the context) if for any $m$ in $M$ there exists $n\in{\mathbb{N}}_{>0}$ such that $nm$ belongs to $M[\underline{S}]$, i.e. if $M/(M[\underline{S}])$ is a torsion abelian group.
@@
    A family $(a_{i}:i\in I)$ of homogeneous elements in $A$ is called $M$-relevant if the multiplicative subset generated by the $a_{i}$’s is relevant.
@@
    for any relevant element $x\in A\otimes_{R}A^{\prime}$, there exists a positive integer $k$, a finite set $J$, and for any $j\in J$ an $M$-relevant element $s_{j}$ in $A^{\prime}$ and an $M^{\prime}$-relevant element $s^{\prime}_{j}$ in $A$ such that $x^{k}=\sum_{j\in J}s_{j}\otimes s^{\prime}_{j}$;
@@
Assume that $M^{\prime}=\{0\}$, so that $A^{\prime}$ is simply an $R$-algebra, which we will denote $R^{\prime}$ for clarity, and we consider the $M$-graded algebra $A\otimes_{R}R^{\prime}$. In this case we have $R^{\prime}_{\dagger}=R^{\prime}$ (see Remark 2.4(2)), hence the proposition says that $\mathrm{Rad}((A\otimes_{R}R^{\prime})_{\dagger})=\mathrm{Rad}(\langle A_{\dagger}\otimes_{R}R^{\prime}\rangle)$. This implies that the preimage of the closed subset $V(A_{\dagger})\subset{\operatorname{Spec}}(A)$ under the projection
@@
Now, assume that $\Psi$ is surjective. Then $\Psi(A_{\dagger})$ is an ideal, and by (1) we have $\mathrm{Rad}(\Psi(A_{\dagger}))\subset{\operatorname{Rad}}(B_{\dagger})$. Let $f\in B$ be a relevant element. Then there exists $N\in\mathbb{Z}_{\geqslant 1}$, nonzero homogeneous elements $f_{1},\dots,f_{r}\in B$ of respective degrees $m_{1},\dots,m_{r}$, and a homogeneous element $g\in B$ such that $f^{N}=f_{1}\cdots f_{r}\cdot g$ and the subgroup of $M$ generated by $m_{1},\dots,m_{r}$ has finite index. Choose homogeneous elements $\tilde{f}_{1},\dots,\tilde{f}_{r},\tilde{g}\in A$ whose respective images are $f_{1},\dots,f_{r},g$. Then $\tilde{f}=\tilde{f}_{1}\cdots\tilde{f}_{r}\cdot\tilde{g}$ is relevant and satisfies $\Psi(\tilde{f})=f^{N}$, hence $f\in\mathrm{Rad}(\Psi(A_{\dagger}))$. This implies that $B_{\dagger}\subset\mathrm{Rad}(\Psi(A_{\dagger}))$, hence $\mathrm{Rad}(B_{\dagger})\subset\mathrm{Rad}(\Psi(A_{\dagger}))$, which finishes the proof of (2). ∎
@@ 2.3 Potions
Let $M$ be an abelian group, and $A$ a commutative $M$-graded ring.
@@
    If $Q$ is a graded $A$-module, we will also denote by $Q_{(S)}$ the degree-00 part $(Q_{S})_{0}$ in $Q_{S}$; it admits a canonical structure of $A_{(S)}$-module.
Note that in the setting of Definition 2.9, by (2.1)–(2.2) we have canonical identifications
$A_{(\underline{S})}\cong A_{(S)},\quad Q_{(\underline{S})}\cong Q_{(S)}.$
If $\{a_{i}:i\in I\}$ is a family of homogeneous elements of $A$, we will denote by
$A_{(\{a_{i}:i\in I\})}$
the potion associated with the multiplicative subset of $A$ generated by $\{a_{i}:i\in I\}$; in case #$\#I=1$ we will write $A_{(a)}$ for $A_{(\{a\})}$.
@@
    Let $f_{1},\ldots,f_{n}\in A$ be relevant homogeneous elements of the same degree. Then we have a canonical open immersion
@@
(3) This follows immediately from (2), since finite localizations of rings induce open immersions.
(4) For $i\in\{1,\ldots,n\}$, put $f_{i}^{\prime}=\frac{f_{i}}{f_{1}+\cdots+f_{n}}\in A_{(f_{1}+\cdots+f_{n})}$ and
@@
where $(A_{(f_{1}+\cdots+f_{n})})_{f_{i}^{\prime}}$ is the localization of $A_{(f_{1}+\cdots+f_{n})}$ with respect to the multiplicative subset generated by $f_{i}^{\prime}$. We have $f_{1}^{\prime}+\cdots+f_{n}^{\prime}=1$ in $A_{(f_{1}+\cdots+f_{n})}$, so by [StP, Tag 01HS] we obtain that
@@
Now by (2) we have $(A_{(f_{1}+\cdots+f_{n})}\big{)}_{f_{i}^{\prime}}=A_{\big{(}(f_{1}+\cdots+f_{n})f_{i}\big{)}}$ and therefore $D(f_{i}^{\prime})\subset{\operatorname{Spec}}(A_{(f_{i})})$, which finishes the proof. ∎
@@ 3.1 Proj scheme of a graded ring
over ${\operatorname{Spec}}(A_{0})$. The $M$-grading on $A$ defines an action of $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$ on ${\operatorname{Spec}}(A)$; we will denote by
@@
In [BS03, Definition 2.2], Brenner and Schröer define the Proj scheme associated with $A$ as a certain open subset in the ringed space $\mathrm{Quot}(A)$ obtained as the cokernel (in the sense of [SGA3, Exp. V, §1.b]) of the maps $a$ and $p$. In Construction 3.1 below we explain an equivalent description of this scheme, obtained by gluing spectra of certain potion rings. (See the comments following the construction for a justification that our description is indeed equivalent with that in [BS03].)
@@
Proof.
This follows from [StP, Tag 01KO], since by construction the intersections in ${\operatorname{Proj}}^{M}(A)$ of affine open subschemes of the form $D_{\dagger}(S)$ are affine. ∎
@@
Using the notation introduced in Definition 2.3(3), in case $S=a^{\mathbb{N}}$ where $a\in A$ is $M$-relevant, we will write $D_{\dagger}(a)$ for $D_{\dagger}(a^{\mathbb{N}})$.
If $f\in A$ is relevant, for any homogeneous $g\in A$ the product $fg$ is relevant.
@@
    If $g$ is also relevant, then we have
    $D_{\dagger}(fg)=D_{\dagger}(f)\cap D_{\dagger}(g).$
@@
    If $\deg(g)=0$, then $D_{\dagger}(fg)$ is the preimage of $D(g)\subset{\operatorname{Spec}}(A_{0})$ under the composition
@@
(2) This follows similarly from Remark 3.2(3).
@@
${\operatorname{Proj}}^{M}(A)=\bigcup_{f\in A\text{ relevant}}D_{\dagger}(f),$ (3.3)
which justifies that the scheme considered in Construction 3.1 coincides with that defined in [BS03, Definition 2.2]. In particular the scheme ${\operatorname{Proj}}^{M}(A)$ can be constructed by gluing open subschemes associated with relevant elements along their intersections, which by Lemma 3.4(2) are again open subschemes associated with relevant elements.
@@
The following result is an immediate corollary of [BS03, Proposition 3.3]. Given a finite homogeneous family $T=(a_{i}:i\in I)$, generating a multiplicative subset $S$ of $A$, we will denote by $C_{T}\subset M\otimes_{\mathbb{Z}}\mathbb{R}$ the closed convex cone $M[\underline{S}\rangle_{\mathbb{R}_{\geqslant 0}}$ (cf. Definition 2.1). Note that $T$ is relevant if and only if $C_{T}$ has nonempty interior.
@@
This is an immediate corollary of [BS03, Proposition 3.3] and the comment preceding (3.3). ∎
@@
In the description given in (3.3), this morphism can be understood as follows. For any $f\in A$ relevant, we have a canonical embedding $A_{(f)}\subset A_{f}$, which provides a morphism ${\operatorname{Spec}}(A_{f})\to{\operatorname{Spec}}(A_{(f)})$. The latter morphisms glue to a morphism from
@@
This morphism is discussed in [BS03, Comments after Definition 2.2]. The closed subscheme $V(A_{\dagger})\subset{\operatorname{Spec}}(A)$ associated with $A_{\dagger}$ is stable under the action of $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$; there is hence also an action on the open complement ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})$. This action stabilizes each open subscheme ${\operatorname{Spec}}(A_{f})$, and by [BS03, Lemma 2.1] the morphism ${\operatorname{Spec}}(A_{S})\to{\operatorname{Spec}}(A_{(f)})$ is a geometric quotient for the action of $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$ in the sense of [MFK93, Definition 0.6]. It follows that (3.4) is also a geometric quotient. As explained in [BS03], this morphism is also affine.
@@
factors through a morphism
$H\times_{{\operatorname{Spec}}(A_{0})}{\operatorname{Proj}}^{M}(A)\to{\operatorname{Proj}}^{M}(A).$
It is easily seen that this morphism defines an action of $H$ on ${\operatorname{Proj}}^{M}(A)$.
Example 3.8.

### v4

@@ 1.3 Contents
We now present the results contained in this document. Let as above $M$ be an abelian group and $A$ be a commutative ring endowed with a grading by $M$. In Sections 2 and 3 we define the notions of $M$-relevant families of $A$ (following Brenner–Schröer, see Definition 2.3) and of potions associated with these families, and explain the definition of the Proj scheme ${\operatorname{Proj}}^{M}(A)$ of $A$ (in case $M$ is finitely generated). The potion associated with a relevant family $(f_{i}:i\in I)$ is the degree-zero part in the localization of $A$ with respect to the multiplicative subset generated by $(f_{i}:i\in I)$; see Definition 2.9. The spectra ${\operatorname{Spec}}(A_{(S)})$ are called potion schemes and are denoted $D_{\dagger}(S)$ (cf. Construction 3.1), and the scheme ${\operatorname{Proj}}^{M}(A)$ is defined by glueing the potions schemes $D_{\dagger}(S)$ over all relevant families (cf. Construction 3.1). This glueing is possible by the magic of potions (cf. Proposition 2.10). In other words, Section 2 focuses on commutative algebra around potions while Section 3 applies this material to explain the construction of Proj schemes.
In Sections 2 and 3, we also prove some basic results that may be known to experts. In particular, using the concept of quasi-relevant element (Definition 2.11), we prove that the Proj scheme of a tensor product of graded rings (with the product grading) is the fiber product of the Proj schemes of the rings (Proposition 3.15). We prove some compatibility results regarding the radical of the ideal generated by relevant elements (Propositions 2.5 and 2.7). In Example 2.13, we explain the relation between potions and dilatations of rings (as studied in [MRR23, Ma23b]). We also show that the Proj construction is functorial (Proposition 3.12). In §3.6 we discuss multi-centered blowups and its relations with dilatations.
In Section 4 we explain how to realize flag varieties of reductive groups, and some vector bundles attached to representations of a Borel subgroup, as the Proj scheme of a natural graded ring (without any choice of strictly dominant character). Finally, in Section 5 we study, given an $M$-graded commutative ring $A$, the natural functors relating the categories of $M$-graded $A$-modules and quasi-coherent sheaves on ${\operatorname{Proj}}^{M}(A)$. In particular, we explain how, under a certain technical assumption satisfied in many examples of interest, one can describe $\mathsf{QCoh}({\operatorname{Proj}}^{M}(A))$ as a Serre quotient of the category of $M$-graded $A$-modules, and therefore obtain a version of Serre’s celebrated theorem.
Many of these results (in particular, those concerned with quasi-coherent sheaves) are counterparts of well-known results regarding Grothendieck’s Proj construction, which can be found e.g. in [StP].
@@ 1.6.1 Rings
All the rings we will consider in the paper will be tacitly assumed to be associative and unital. (In practice, all the rings we want to consider will be commutative, but we will mention this assumption when it is necessary.)
@@ 1.6.2 Graded rings
Let us make explicit our conventions on graded rings. For this we fix an arbitrary abelian group $M$. An $M$-graded ring is a ring $A$ endowed with a direct sum decomposition $A=\bigoplus_{m\in M}A_{m}$ such that $A_{m}\cdot A_{m^{\prime}}\subset A_{m+m^{\prime}}$ for all $m,m^{\prime}\in M$.
Let $A$ be an $M$-graded ring. An element $a\in A$ is called homogeneous if it belongs to $A_{m}$ for some $m\in M$. In this case, if $a$ is nonzero, then the element $m$ is unique; it is denoted $\deg(a)$ and called the degree of $a$.
@@ 1.6.3 Other conventions
If $A$ is a ring and $a\in A$ is an element, we set $a^{\mathbb{N}}:=\{a^{n}:n\in{\mathbb{N}}\}\subset A$. (Here ${\mathbb{N}}$ denotes the set of non negative integers; this does not follow the conventions of [StP].) For any $a\in A$, by convention we have $a^{0}=1$.
@@
If $A$ is a commutative ring and $I\subset A$ is an ideal, we will denote by $V(I)\subset{\operatorname{Spec}}(A)$ the closed subset defined by $I$. If $f\in A$ we will denote by $D(f)\subset{\operatorname{Spec}}(A)$ the open subscheme defined by $f$, i.e. the complement of $V(A\cdot f)$.
If $N$ is any commutative monoid, there is a universal homomorphism from $N$ to a group $N^{\mathrm{gp}}$ (cf. [Og18, §1.2]). In case $N$ is cancellative [^1] The terminology used in [Og18] is integral. (i.e. if the equality $x+y=x^{\prime}+y$ implies that $x=x^{\prime}$ for any $x,x^{\prime},y\in N$) this map is injective. If $N$ is a submonoid of a given abelian group $M$, then $N$ is cancellative and the group $N^{\mathrm{gp}}$ naturally identifies with the subgroup $N-N:=\{n-n^{\prime}:n,n^{\prime}\in N\}$ of $M$.
@@ 2.2 Relevant families
Let $S$ be a homogeneous multiplicative subset of $A$. We denote by $\deg(S)$ the subset of $M$ consisting of degrees of nonzero homogeneous elements of $S$.
@@
    A homogeneous multiplicative subset $S$ of $A$ is called $M$-relevant (or just relevant if $M$ is clear from the context) if for any $m$ in $M$ there exists $n\in{\mathbb{Z}}_{>0}$ such that $nm$ belongs to $M[\underline{S}]$, i.e. if $M/(M[\underline{S}])$ is a torsion abelian group.
@@
    A family $\{a_{i}:i\in I\}$ of homogeneous elements in $A$ is called $M$-relevant if the multiplicative subset generated by the $a_{i}$’s is relevant.
@@
    for any relevant element $x\in A\otimes_{R}A^{\prime}$, there exists a positive integer $k$, a finite set $J$, and for any $j\in J$ an $M$-relevant element $s_{j}$ in $A$ and an $M^{\prime}$-relevant element $s^{\prime}_{j}$ in $A^{\prime}$ such that $x^{k}=\sum_{j\in J}s_{j}\otimes s^{\prime}_{j}$;
@@
Assume that $M^{\prime}=\{0\}$, so that $A^{\prime}$ is simply an $R$-algebra, which we will denote $R^{\prime}$ for clarity, and we consider the $M$-graded algebra $A\otimes_{R}R^{\prime}$. In this case we have $R^{\prime}_{\dagger}=R^{\prime}$ (see Remark 2.4(2)), hence the proposition says that $\mathrm{Rad}((A\otimes_{R}R^{\prime})_{\dagger})=\mathrm{Rad}(\langle A_{\dagger}\otimes_{R}R^{\prime}\rangle)$, which is exactly [BS07, Lemma 2.7]. This statement implies that the preimage of the closed subset $V(A_{\dagger})\subset{\operatorname{Spec}}(A)$ under the projection
@@
Now, assume that $\Psi$ is surjective. Then $\Psi(A_{\dagger})$ is an ideal, and by (1) we have $\mathrm{Rad}(\Psi(A_{\dagger}))\subset{\operatorname{Rad}}(B_{\dagger})$. Let $f\in B$ be a relevant element. Then there exists $N\in\mathbb{Z}_{\geqslant 1}$, elements $m_{1},\dots,m_{r}\in M$ which generate a subgroup of finite index, and for any $i$ an element $f_{i}\in B_{m_{i}}$, such that $f^{N}=f_{1}\cdots f_{r}$. Choose for any $i$ a preimage $\tilde{f}_{i}\in A_{m_{i}}$ for $f_{i}$. Then $\tilde{f}=\tilde{f}_{1}\cdots\tilde{f}_{r}$ is relevant and satisfies $\Psi(\tilde{f})=f^{N}$, hence $f\in\mathrm{Rad}(\Psi(A_{\dagger}))$. This implies that $B_{\dagger}\subset\mathrm{Rad}(\Psi(A_{\dagger}))$, hence $\mathrm{Rad}(B_{\dagger})\subset\mathrm{Rad}(\Psi(A_{\dagger}))$, which finishes the proof of (2). ∎
@@ 2.3 Potions
Let $M$ be an abelian group, and let $A$ be a commutative $M$-graded ring.
@@
    If $Q$ is a graded $A$-module, we will also denote by $Q_{(S)}$ the degree-00 part $(Q_{S})_{0}$ of $Q_{S}$; it admits a canonical structure of $A_{(S)}$-module.
Note that in the setting of Definition 2.9, by (2.1)–(2.2) we have canonical identifications $A_{(\underline{S})}\cong A_{(S)}$ and $Q_{(\underline{S})}\cong Q_{(S)}$. If $\{a_{i}:i\in I\}$ is a family of homogeneous elements of $A$, we will denote by $A_{(\{a_{i}:i\in I\})}$ the potion associated with the multiplicative subset of $A$ generated by $\{a_{i}:i\in I\}$; in case #$\#I=1$ we will write $A_{(a)}$ for $A_{(\{a\})}$.
@@
    Let $f_{1},\ldots,f_{n}\in A$ be nonzero relevant homogeneous elements of the same degree. Then we have a canonical open immersion
@@
(3) In case $T$ contains 00, the claim is clear. Otherwise, it follows immediately from (2), since finite localizations of rings induce open immersions.
(4) For $i\in\{1,\ldots,n\}$, put $f_{i}^{\prime}=\frac{f_{i}}{f_{1}+\cdots+f_{n}}\in A_{(f_{1}+\cdots+f_{n})}$ and consider the open subscheme
@@
where $(A_{(f_{1}+\cdots+f_{n})})_{f_{i}^{\prime}}$ is the localization of $A_{(f_{1}+\cdots+f_{n})}$ with respect to the multiplicative subset generated by $f_{i}^{\prime}$. We have $f_{1}^{\prime}+\cdots+f_{n}^{\prime}=1$ in $A_{(f_{1}+\cdots+f_{n})}$, so by Iten(3) in [StP, Tag 01HS] we obtain that
@@
Now by (2) we have $(A_{(f_{1}+\cdots+f_{n})}\big{)}_{f_{i}^{\prime}}=A_{\big{(}(f_{1}+\cdots+f_{n})f_{i}\big{)}}$ and therefore $D(f_{i}^{\prime})$ identifies with an open subscheme in ${\operatorname{Spec}}(A_{(f_{i})})$, which finishes the proof. ∎
@@ 3.1 Proj scheme of a graded ring
over ${\operatorname{Spec}}(A_{0})$ associated with $M$. The $M$-grading on $A$ defines an action of $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$ on ${\operatorname{Spec}}(A)$; we will denote by
@@
In [BS03, Definition 2.2], Brenner and Schröer define the Proj scheme [^3] Brenner–Schröer use the terminology “homogeneous spectrum” of $A$. We prefer to use the term “Proj scheme” to emphasize the parallel with Grothendieck’s construction. associated with $A$ as a certain open subset in the ringed space $\mathrm{Quot}(A)$ obtained as the cokernel (in the sense of [SGA3, Exp. V, §1.b]) of the maps $a$ and $p$. In Construction 3.1 below we explain an equivalent description of this scheme, obtained by gluing spectra of certain potion rings. (See §3.2 for a justification that our description is indeed equivalent to that in [BS03].)
@@
Example 3.3.
If $\{a_{i}:i\in I\}$ is a finite relevant family of elements of $A$, we will write $D_{\dagger}(\{a_{i}:i\in I\})$ for the open subscheme defined by the multiplicative subset of $A$ generated by the $a_{i}$’s. In case #$\#I=1$ we will write $D_{\dagger}(a)$ for $D_{\dagger}(\{a\})$.
3.2 First properties
By construction, ${\operatorname{Proj}}^{M}(A)$ is covered by the affine open subschemes $D_{\dagger}(S)$, which have the property that the intersection of any two such subschemes is affine. This implies that the diagonal morphism of ${\operatorname{Proj}}^{M}(A)$ is affine (as in [BS03, Proposition 3.1]), hence (since affine morphisms are quasi-compact) the following statement.
@@ 3.2 First properties
For the next statement, note that if $f\in A$ is relevant, for any homogeneous $g\in A$ the product $fg$ is relevant.
@@
    If $g$ is also relevant, then we have $D_{\dagger}(fg)=D_{\dagger}(f)\cap D_{\dagger}(g)$.
@@
    If $g\in A_{0}$, then $D_{\dagger}(fg)$ is the preimage of $D(g)\subset{\operatorname{Spec}}(A_{0})$ under the composition
@@
(2) This follows similarly from Remark 3.2(3) since $(fg)^{\mathbb{N}}\subset f^{\mathbb{N}}\cdot g^{\mathbb{N}}\subset\underline{(fg)^{\mathbb{N}}}$.
@@
${\operatorname{Proj}}^{M}(A)=\bigcup_{f\in A\text{ relevant}}D_{\dagger}(f).$ (3.3)
In fact, in view of Remark 3.2(2) we can restrict this union to non nilpotent relevant elements.
Comparing Lemma 3.5(2) with the proof of [BS03, Proposition 3.1] we see that the intersections of these open subschemes are as in the scheme constructed in [BS03], which justifies that our scheme ${\operatorname{Proj}}^{M}(A)$ coincides with that defined in [BS03, Definition 2.2]. In particular the scheme ${\operatorname{Proj}}^{M}(A)$ can be constructed by gluing open subschemes associated with relevant elements along their intersections, which are again open subschemes associated with relevant elements. We however feel that the possibility of defining open subschemes associated with relevant families adds some flexibility which might be useful.
@@
The following result is an immediate corollary of [BS03, Proposition 3.3]. Given a finite family $T=\{a_{i}:i\in I\}$ of homogeneous elements of $A$, generating a multiplicative subset $S$, we will denote by $C_{T}\subset M\otimes_{\mathbb{Z}}\mathbb{R}$ the closed convex cone $M[\underline{S}\rangle_{\mathbb{R}_{\geqslant 0}}$ (cf. Definition 2.1). Note that $T$ is relevant if and only if $C_{T}$ has nonempty interior.
@@
This follows from [BS03, Proposition 3.3] and the comment preceding (3.3). ∎
@@
In the description given in (3.3), this morphism can be understood as follows. For any $f\in A$ relevant, we have a canonical embedding $A_{(f)}\subset A_{f}$, which provides a morphism ${\operatorname{Spec}}(A_{f})\to{\operatorname{Spec}}(A_{(f)})$. The latter morphisms glue to the desired morphism from
@@
This morphism is affine since the preimage of ${\operatorname{Spec}}(A_{(f)})$ is ${\operatorname{Spec}}(A_{f})$, for any relevant $f$.
The morphism (3.4) is discussed further in [BS03, Comments after Definition 2.2]. The closed subscheme $V(A_{\dagger})\subset{\operatorname{Spec}}(A)$ associated with $A_{\dagger}$ is stable under the action of $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$; there is hence also an action on the open complement ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})$. This action stabilizes each open subscheme ${\operatorname{Spec}}(A_{f})$, and by [BS03, Lemma 2.1] the morphism ${\operatorname{Spec}}(A_{f})\to{\operatorname{Spec}}(A_{(f)})$ is a geometric quotient for the action of $\mathrm{D}_{{\operatorname{Spec}}(A_{0})}(M)$ in the sense of [MFK93, Definition 0.6]. It follows that (3.4) is also a geometric quotient.
@@
factors through a morphism $H\times_{{\operatorname{Spec}}(A_{0})}{\operatorname{Proj}}^{M}(A)\to{\operatorname{Proj}}^{M}(A)$. It is easily seen that this morphism defines an action of $H$ on ${\operatorname{Proj}}^{M}(A)$.